# Machine-precision energy conservative quadrature hyperreduction of Lagrangian hydrodynamics

### v1

@@ 4.1 Energy conservation
The semidiscrete full model (2.3) conserves total energy by design: $\frac{d}{dt}TE(\bm{w}(t))=0$ for all $t\geq 0$. In addition, the RK2A discrete scheme (2.8) conserves the discrete total energy: $TE(\bm{w}_{k+1})=TE(\bm{w}_{k})$ for all $k\geq 0$ [72]. In the following theorem we derive sufficient conditions that ensure that the conservation of discrete total energy is preserved in the discrete hyperreduced model (4.2).
@@
Then the discrete hyperreduced model (4.2) conserves the discrete total energy
@@
The three conditions presented in Theorem 1 are motivated by the physics and structure of the hyperreduced model. The first two conditions ensure that the kinetic energy $KE$ and the internal energy $IE$ respectively can be represented in terms of the reduced state vectors. The third condition is the main structure preserving property and couples the velocity and energy forcing terms, motivated by the fact that the fluid flow governed by the discrete hyperreduced model (4.2) is driven by pressure-volume work.
@@ 4.2 Implementation
To implement the energy conservative EQP method we begin with the semidiscrete model (3.7). We use zero offset vectors for position, velocity and energy in (3.1), which ensures we satisfy condition 1 of Theorem 1. To meet condition 2 we enrich the energy basis matrix $\bm{\Phi}_{e}$ by adding the energy unit $\bm{1}_{\mathcal{E}}$ as a column vector. Moreover, we orthonormalize the basis matrices $\bm{\Phi}_{v}$ and $\bm{\Phi}_{e}$ such that the reduced mass matrices (3.8) reduce to identity matrices. To satisfy condition 3 we derive one combined reduced quadrature rule to be used for both the velocity $\bm{F}_{v}^{\phi}$ and energy $\bm{F}_{e}^{\phi}$ force functions. More specifically, we form a reduced quadrature rule for the force function $\bm{F}^{\phi}:\mathbb{R}^{N}\to\mathbb{R}^{n_{v}\times n_{e}}$
@@
where $\bm{G}$ denotes the integrand taking values in $\mathbb{R}^{n_{v}\times n_{e}}$ with entries
@@
Using $\bm{F}^{\phi}$ the velocity and energy force functions are given by
@@
Starting with the full quadrature rule we employ the EQP method to derive a reduced quadrature rule $\{x_{j},\tilde{\rho}_{j}\}_{j=0}^{J-1}$ with $\hat{J}\ll J$ nonzero weights. We use the reduced rule to define the reduced force function $\hat{\bm{F}}^{\phi}:\mathbb{R}^{n}\to\mathbb{R}^{n_{v}\times n_{e}}$
@@
that approximates the evaluation of $\bm{F}_{v}^{\phi}$ and $\bm{F}_{e}^{\phi}$ with reduced computational cost
@@
satisfying condition 3 of Theorem 1. Therefore, we conclude that the discrete hyperreduced model (4.2) conserves the discrete total energy for all time indices $k\geq 0$. Additional details on the implementation of the energy conservative EQP method are given in A.
@@ 5.1 Problems
In the following paragraphs we give an outline of each problem case and its simulation parameters. The spatial domain in each problem is discretized using an initially uniform, cartesian mesh of hexahedral elements. Table 1 presents discretization parameters used for each of the simulated problems, including the number of basis functions used for each finite element space and the corresponding polynomial bases orders.
@@
Figure 1: Final snapshots of the velocity field for each simulated problem case. The top row corresponds to the results obtained using the full model; the middle row using the model hyperreduced by the basic EQP method; the bottom row using the model hyperreduced by the energy conservative EQP method. From left to right, the first column corresponds to the Sedov problem; the second to Gresho; the third to triple point; the fourth to Taylor-Green.
@@
The computational domain is $\Omega_{0}=[0,1]^{3}$. We use a constant adiabatic index $\gamma=1.4$, initial density $\rho=1$, and zero initial velocity. The initial energy has magnitude $|e|=0.25$ and is all concentrated at the domain origin. We employ one of the artificial viscosity stress tensors developed in [72] which preserves the propagating shock’s radial symmetry. The simulation time interval is $[0,0.3]$.
@@
$p(r,\phi)=\begin{cases}5+12.5r^{2}&\text{if $0\leq r<0.2$,}\\ 9-4\log(0.2)+12.5-20r+4\log r&\text{if $0.2\leq r<0.4$,}\\ 3+4\log 2&\text{if $r\geq 0.4$,}\end{cases}$
@@
$v_{\phi}(r,\phi)=\begin{cases}5r&\text{if $0\leq r<0.2$,}\\ 2-5r&\text{if $0.2\leq r<0.4$,}\\ 0&\text{if $r\geq 0.4$.}\end{cases}$
@@
$\gamma=\begin{cases}1.5&\text{if $x\leq 1$ or $y>1.5$,}\\ 1.4&\text{if $x>1$ and $y\leq 1.5$,}\end{cases}$
@@
$\rho=\begin{cases}1&\text{if $x\leq 1$ or $y\leq 1.5$,}\\ 1/8&\text{if $x>1$ and $y>1.5$,}\end{cases}$
@@
$p=\begin{cases}1&\text{if $x\leq 1$,}\\ 0.1&\text{if $x>1$.}\end{cases}$
The initial energy is computed from the thermodynamic equation of state. Given the formation of shock waves, we also use an artificial viscosity tensor [72]. The simulated time interval is $[0,0.8]$.
@@
The initial energy is computed using the thermodynamic equation of state. We use no artificial viscosity stress. The simulated time interval is $[0,0.4]$.
@@ 5.2 Results
Figure 1 presents the final velocity snapshots obtained using the full model and the two reduced models for each problem case. Table 2 presents the numerical results for the four problem cases using the two EQP hyperreduction methods compared in this work.
@@
In terms of the differences in relative errors achieved for the position, velocity and energy solutions, the two methods perform similarly in Sedov and triple point but demonstrate more pronounced discrepancies in Gresho and Taylor-Green. More specifically, for the Sedov case both methods lead to similar errors with CEQP outperforming BEQP in position and velocity and conversely in energy. Similarly, the two methods yield similar results for triple point with BEQP performing better by about an order of magnitude in all three variables. The situation is different for Taylor-Green and Gresho. For Taylor-Green CEQP underperforms by about 2-3 orders of magnitude in the three solution errors, while for Gresho by about 3-4 orders of magnitude. Despite the more pronounced differences in accuracy in these problem cases, the errors obtained by CEQP are still reasonably low.
@@ 5.3 Discussion
Among the presented numerical results, the discrepancy in solution accuracy between the basic and energy conservative EQP methods is largest in the Gresho problem case. In addition, Gresho is the problem case generating the largest values of solution errors for CEQP, despite the fact that its energy conservation performance is on par with the other experiments. As of this writing, it is not well understood why that is the case, and whether the elevated errors are due to the numerical implementation of the method or due to some intrinsic challenge posed by the Gresho problem case. We intend to continue investigating this issue.
@@
The speedup factors reported in the numerical results are relatively small compared to [33] for example. This is partly because the results were generated using a numerical implementation of the basic and energy conservative EQP methods that is not fully optimized yet. More specifically, the used implementation still performs some of the lifting of the reduced state vectors to full state vectors using all degrees of freedom of the system, instead of using only the degrees of freedom sampled by the reduced quadrature rules. Independently of those optimizations, we also expect that the speedup values will increase with increasing values for the mesh refinement parameter $m$ and the FEM basis polynomial order $k$.
@@
A direction of future research is comparing the performance of different NNLS solver algorithms, especially as it relates to the sparsity and accuracy of the derived reduced quadrature rules. As shown in Table 2 the Lawson-Hanson algorithm used in this work leads to sparse quadrature rules. Nevertheless, it is not clear how different methods would compare against our reported results. In addition, we intend to investigate why the $LQ$ preconditioning of the NNLS constraints matrix (ensuring orthogonality of constraints) leads to improved performance of the Lawson-Hanson algorithm. Although observed empirically, to our knowledge it is not currently understood why that happens [80].
@@
In the present work we have focused on reproductive simulations, where each reduced model is tested using initial conditions that have been used in its training. The methods studied in this work readily generalize to predictive simulations, where out of sample initial conditions are used to test the model’s predictive performance. Using the methods developed in this work for predictive simulations is something we intend to pursue in the future.
@@
The approach we employ in this work to derive reduced models conceptually builds upon the finite element framework. Starting with the FEM discretization of the problem using the FEM basis functions as the trial and test functions, we proceed to replace these functions by ones that are derived in a data driven fashion from simulation data. One important difference between the reduced models derived in this way and the original FEM model is that the latter employs trial and test functions that are local in space, meaning that their support is confined within a small number of elements. On the contrary, the reduced models employ trial and/or test functions with a support that is generally extended across a large number of elements, possibly including the whole spatial domain. Although this allows one to use a smaller number of degrees of freedom to represent the considered dynamics, it can also lead to some of the extrapolation problems encountered by reduction methods to date, since the accurate representation and prediction of out of sample dynamics cannot be guaranteed.
@@
There have been attempts to derive data driven model reduction methods that use spatially local reduced basis functions, allowing them to further explore potential connections with the finite element framework [81, 82, 83, 84, 85]. It is interesting to explore the applicability of such methods to the problem setting considered in the present work, especially since quadrature hyperreduction methods are naturally well suited to the finite element framework.
@@ 6 Conclusion
We have developed reduced models of nonlinear Lagrangian hydrodynamics using quadrature hyperreduction methods to approximate the evaluation of the nonlinear terms with reduced computational cost. The studied hyperreduction methods are based on the empirical quadrature procedure (EQP) which is well suited to the finite element discretization of the studied nonlinear conservation laws.
Starting with an application of the basic EQP method, we have developed a variant of the method that preserves the energy conservation of the full model during the reduction process. In addition to proving exact energy conservation in the resulting reduced model, we have also performed numerical simulations of four different benchmark problems demonstrating the conservation of energy to near machine precision. The additional benefit of energy conservation has been obtained with only a moderate reduction in the accuracy or speedup of the reduced simulations compared to the basic EQP method.
@@ A.1 Basic EQP
In this section we offer details on the nonnegative least squares (NNLS) formulation used for the basic EQP method developed in Section 3.1, as well as on preconditioning operations performed before soling the formed NNLS problem.
@@
To solve the linear optimization problem for the reduced quadrature rule described in Section 3.1 we reformulate it as a nonnegative least squares problem for the vector of quadrature weights $\tilde{\bm{\rho}}^{v}\in\mathbb{R}^{J_{v}}$ [59, 79]. Given the known vector of full quadrature weights $\bm{\rho}^{v}\in\mathbb{R}^{J_{v}}$ we form the accuracy constraints matrix $\bm{C}_{v}\in\mathbb{R}^{N_{c}^{v}\times J_{v}}$, $N_{c}^{v}\ll J_{v}$ with entries
@@
with $0\leq s=i+kn_{v}<N_{c}^{v}$ and the vector $\bm{b}=\bm{C}_{v}\bm{\rho}^{v}\in\mathbb{R}^{N_{c}}$. The vector $\tilde{\bm{\rho}}^{v}$ can now be found as the solution to the nonnegative least squares problem
@@
More specifically, for a given vector of error thresholds $\bm{\epsilon}\in\mathbb{R}^{N_{c}}$ we use the iterative Lawson-Hanson algorithm to solve the following problem for $\bm{r}\geq\bm{0}$
@@
where $\bm{c}_{v}^{s}\in\mathbb{R}^{J_{v}}$ denotes the $s$th row of matrix $\bm{C}_{v}$ and $b_{s}$, $\epsilon_{s}$ the $s$th entries of vectors $\bm{b}$ and $\bm{\epsilon}$ respectively [86, 87].
In the NNLS formulation equation (A.1) is used to enforce the accuracy constraints of the linear optimization problem. The nonnegativity constraints are enforced implicitly by the employed algorithm. Finally, compared to the original form of the problem, the NNLS formulation does not explicitly promote sparsity by minimizing the $\ell^{1}$ norm of solution vector $\tilde{\bm{\rho}}^{v}$. Nevertheless, several numerical experiments (including our own) show that significantly sparse solutions are obtained when the Lawson-Hanson algorithm is employed [59, 79, 78].
@@
Before solving equation (A.1) we perform two conditioning operations on constraints matrix $\bm{C}_{v}$ to reduce the number of iterations required by the Lawson-Hanson algorithm to converge and to increase the derived solution’s accuracy [80]. First, we rescale each row of $\bm{C}_{v}$ so that its entries have absolute value less than or equal to one. Second, we use the reduced LQ decomposition of $\bm{C}_{v}=\bm{L}_{v}\bm{Q}_{v}$, where $\bm{L}_{v}\in\mathbb{R}^{N_{c}^{v}\times N_{c}^{v}}$ is lower triangular and $\bm{Q}_{v}\in\mathbb{R}^{N_{c}^{v}\times J_{v}}$ has orthonormal rows, to transform the NNLS problem (A.1) to
@@
where $\tilde{\bm{b}}=\bm{Q}_{v}\bm{\rho}^{v}\in\mathbb{R}^{N_{c}^{v}}$ with the assumption that $\bm{L}_{v}$ is nonsingular. In addition, we modify the error thresholds $\epsilon_{s}$ accordingly to ensure that the solution obtained from (A.3) will also satisfy (A.2) with the original error parameters. To that end we define the new error thresholds $\tilde{\epsilon}_{s}$
@@
which are sufficient to ensure that the transformed problem’s solution satisfies (A.2) [80]. The transformed NNLS problem reads

### v2

@@ 4.1 Energy conservation
The semidiscrete full model (2.3) conserves the total energy by design: $\frac{d}{dt}TE(\bm{w}(t))=0$ for all $t\geq 0$. In addition, the RK2A discrete scheme (2.8) conserves the discrete total energy: $TE(\bm{w}_{k+1})=TE(\bm{w}_{k})$ for all $k\geq 0$ [57]. In the following theorem we derive sufficient conditions that ensure that the conservation of discrete total energy is preserved in the discrete hyper-reduced model (4.2).
@@
Then the discrete hyper-reduced model (4.2) conserves the discrete total energy; namely,
@@
The three conditions presented in Theorem 1 are motivated by the physics and structure of the hyper-reduced model. The first two conditions ensure that the kinetic energy $KE$ and internal energy $IE$ respectively can be represented in terms of the reduced state vectors. The third condition is the main structure preserving property coupling the velocity and energy forcing terms, motivated by the fact that the fluid flow governed by the discrete hyper-reduced model (4.2) is driven by pressure-volume work.
@@ 4.2 Implementation
To implement the energy conservative EQP method we begin with the semidiscrete model (3.7). We use zero offset vectors for position, velocity and energy in (3.1), which ensures we satisfy condition 1 of Theorem 1. To meet condition 2 we enrich the energy basis matrix $\bm{\Phi}_{e}$ by adding the energy unit $\bm{1}_{\mathcal{E}}$ as a column vector. Moreover, we orthonormalize the basis matrices $\bm{\Phi}_{v}$ and $\bm{\Phi}_{e}$ such that the reduced mass matrices (3.8) reduce to identity matrices.
To satisfy condition 3 we are going to derive one combined reduced quadrature rule to be used for both the velocity $\bm{F}_{v}^{\phi}$ and energy $\bm{F}_{e}^{\phi}$ force functions. More specifically, we are going to form a reduced quadrature rule for the force function $\bm{F}^{\phi}:\mathbb{R}^{N}\to\mathbb{R}^{n_{v}\times n_{e}}$
@@
where $\bm{G}$ denotes the integrand taking values in $\mathbb{R}^{n_{v}\times n_{e}}$, with entries
@@
With respect to $\bm{F}^{\phi}$, the velocity and energy force functions are given by
@@
Starting with this rule, we employ the EQP method to derive a reduced quadrature rule $\{x_{j},\tilde{\rho}_{j}\}_{j=0}^{J-1}$ with $\hat{J}\ll J$ nonzero weights, and use it to define the reduced force function $\hat{\bm{F}}^{\phi}:\mathbb{R}^{n}\to\mathbb{R}^{n_{v}\times n_{e}}$
@@
The evaluation of $\bm{F}_{v}^{\phi}$ and $\bm{F}_{e}^{\phi}$ is then approximated with reduced computational cost by
@@
satisfying condition 3 of Theorem 1. Therefore, we conclude that the discrete hyper-reduced model (4.2) conserves the discrete total energy for all time indices $k\geq 0$. Additional details on the implementation of the energy conservative EQP method are given in A.
@@ 5.1 Problems
In the following paragraphs we provide an overview of each problem case and its simulation parameters. The spatial domain in each problem is discretized using an initially uniform, cartesian mesh of hexahedral elements. Table 1 presents discretization parameters used for each of the simulated problems, including the number of basis functions used for each finite element space and the corresponding polynomial basis orders.
@@
Figure 1: Final snapshots of the velocity field for each simulated problem case. The top row corresponds to the results obtained using the full model; the middle row using the model hyper-reduced by the basic EQP method; the bottom row using the model hyper-reduced by the energy conservative EQP method. From left to right, the first column corresponds to the Sedov problem; the second to Gresho; the third to triple point; the fourth to Taylor-Green.
@@
The computational domain is $\Omega_{0}=[0,1]^{3}$. We use a constant adiabatic index $\gamma=1.4$, initial density $\rho=1$, and zero initial velocity. The initial energy has magnitude $|e|=0.25$ and is all concentrated at the domain origin. We employ one of the artificial viscosity stress tensors developed in [57] that preserves the propagating shock’s radial symmetry. The simulation time interval is $[0,0.3]$.
@@
$p(r,\phi)=\begin{cases}5+12.5r^{2}&\text{if $0\leq r<0.2$}\\ 9-4\log(0.2)+12.5-20r+4\log r&\text{if $0.2\leq r<0.4$}\\ 3+4\log 2&\text{if $r\geq 0.4$}\end{cases}$
@@
$v_{\phi}(r,\phi)=\begin{cases}5r&\text{if $0\leq r<0.2$}\\ 2-5r&\text{if $0.2\leq r<0.4$}\\ 0&\text{if $r\geq 0.4$.}\end{cases}$
@@
$\gamma=\begin{cases}1.5&\text{if $x\leq 1$ or $y>1.5$}\\ 1.4&\text{if $x>1$ and $y\leq 1.5$}\end{cases}$
@@
$\rho=\begin{cases}1&\text{if $x\leq 1$ or $y\leq 1.5$}\\ 1/8&\text{if $x>1$ and $y>1.5$}\end{cases}$
@@
$p=\begin{cases}1&\text{if $x\leq 1$}\\ 0.1&\text{if $x>1$.}\end{cases}$
The initial energy is computed from the thermodynamic equation of state (2.2). Given the formation of shock waves, we also use an artificial viscosity tensor [57]. The simulated time interval is $[0,0.8]$.
@@
The initial energy is computed using the thermodynamic equation of state (2.2). We use no artificial viscosity stress. The simulated time interval is $[0,0.4]$.
@@ 5.2 Results
Figure 1 presents the final velocity snapshots obtained using the full model and the two reduced models for each problem case. Table 2 presents the numerical results for the four problem cases using the two EQP hyper-reduction methods compared in this work.
@@
In terms of the differences in relative errors achieved for the position, velocity and energy solutions, the two methods perform similarly in Sedov and triple point but demonstrate more pronounced discrepancies in Gresho and Taylor-Green. More specifically, for the Sedov case both methods lead to similar errors, with CEQP outperforming BEQP in position and velocity and conversely in energy. Similarly, the two methods yield similar results for triple point, with BEQP performing better by about an order of magnitude in all three variables. The situation is different for Taylor-Green and Gresho. For Taylor-Green, CEQP underperforms by about 2-3 orders of magnitude in the three solution errors, while for Gresho by about 3-4 orders of magnitude. Despite the more pronounced differences in accuracy in these problem cases, the errors obtained by CEQP are still reasonably low.
@@ 5.3 Discussion
Among the presented numerical results, the discrepancy in solution accuracy between the basic and energy conservative EQP methods is largest in the Gresho problem case. In addition, Gresho is the problem case generating the largest values of solution errors for CEQP, despite the fact that its energy conservation performance is on par with the other experiments. As of this writing, it is not well understood why that is the case, and whether the elevated errors are due to the numerical implementation of the method or due to some intrinsic challenge posed by the Gresho problem.
@@
The reported speedup factors are relatively small compared to those reported in [22], where interpolation methods are used for hyper-reduction. This is partly because our results are generated using a numerical implementation of the basic and energy conservative EQP methods that is not fully optimized yet. More specifically, the used implementation performs some of the lifting of the reduced state vectors to full state vectors using all degrees of freedom, instead of using only the degrees of freedom sampled by the reduced quadrature rules. Independently of those optimizations, we also expect that the speedup values will increase with increasing values for the mesh refinement parameter $m$ and the FEM basis polynomial order $k$.
@@
A direction of future research is comparing the performance of different NNLS algorithms, especially as it relates to the sparsity and accuracy of the derived reduced quadrature rules. As shown in Table 2, the Lawson-Hanson algorithm used in this work leads to sparse quadrature rules. Nevertheless, it is not clear how different methods would compare against our reported results. In addition, we intend to investigate why the $LQ$ preconditioning of the NNLS constraints matrix (ensuring orthogonality of constraints) leads to improved performance of the Lawson-Hanson algorithm. Although observed empirically, to our knowledge it is not fully understood why that happens [63].
@@
In the present work we focus on reproductive simulations, where each reduced model is tested using initial conditions that have been used in its training. The developed methods readily generalize to predictive simulations, where out of sample initial conditions are used to test the model’s performance. We intend to pursue this in the near future.
@@
Conceptually, the model reduction approach employed in this work builds upon the finite element framework. Starting with the FEM discretization of the problem, with the FEM basis functions acting as the trial and test functions, we proceed to replace these functions by ones derived from simulation data. One important difference between the two model formulations is that the FEM model employs trial and test functions that are local in space, meaning that their support is confined within a small number of elements. On the contrary, the reduced models employ trial and/or test functions with a support that is generally extended across a large number of elements, possibly including the whole spatial domain. Although this allows us to use a smaller number of degrees of freedom to represent the considered dynamics, it can also lead to some of the extrapolation problems encountered by reduction methods to date, since the accurate representation and prediction of out of sample dynamics cannot be guaranteed.
@@
There is ongoing work related to deriving data based reduction methods that use spatially local reduced basis functions, bringing them closer to a data based analog of the finite element framework [64, 65, 66, 67, 68]. It is interesting to explore the applicability of such methods to the problem setting considered in the present work, especially since quadrature hyper-reduction methods are naturally well suited to the finite element framework.
@@ 6 Conclusion
We developed reduced models for nonlinear Lagrangian hydrodynamics problems using quadrature hyper-reduction to approximate the evaluation of the nonlinear terms with reduced computational cost. The hyper-reduction was performed using the EQP method, which is well suited to the finite element discretization of the considered nonlinear conservation laws.
Starting with an application of the basic EQP method, we developed a variant of the method that preserves the energy conservation of the full model in the reduction process. In addition to proving exact energy conservation in the resulting reduced model, we performed numerical simulations of four different benchmark problems demonstrating the conservation of energy to near machine precision. Our reported results show that the additional benefit of energy conservation is obtained with only a moderate reduction in the accuracy or speedup of the reduced simulations compared to the basic EQP method.
@@ A.1 Basic EQP
We consider the nonnegative least squares (NNLS) formulation used for the basic EQP method developed in Section 3.1, as well as the preconditioning operations performed before solving the formed NNLS problem.
@@
To solve the linear optimization problem for the reduced quadrature rule described in Section 3.1, we reformulate it as an NNLS problem for the vector of quadrature weights $\tilde{\bm{\rho}}^{v}\in\mathbb{R}^{J_{v}}$ [52, 62]. Given the known vector of full quadrature weights $\bm{\rho}^{v}\in\mathbb{R}^{J_{v}}$, we form the accuracy constraints matrix $\bm{C}_{v}\in\mathbb{R}^{N_{c}^{v}\times J_{v}}$, $N_{c}^{v}\ll J_{v}$, with entries
@@
with row index $0\leq s=i+kn_{v}<N_{c}^{v}$. Denoting $\bm{b}=\bm{C}_{v}\bm{\rho}^{v}\in\mathbb{R}^{N_{c}^{v}}$, the vector $\tilde{\bm{\rho}}^{v}$ is found by solving the NNLS problem
@@
More specifically, for a given vector of error thresholds $\bm{\epsilon}\in\mathbb{R}^{N_{c}}$, we use the iterative Lawson-Hanson algorithm to solve the following problem for $\bm{r}\geq\bm{0}$
@@
where $\bm{c}_{v}^{s}\in\mathbb{R}^{J_{v}}$ denotes the $s$th row of matrix $\bm{C}_{v}$, and $b_{s}$, $\epsilon_{s}$ the $s$th respective entries of vectors $\bm{b}$ and $\bm{\epsilon}$ [69, 70].
In the NNLS formulation, (A.1) is used to enforce the accuracy constraints of the linear optimization problem. The nonnegativity constraints are enforced implicitly by the employed algorithm. Finally, compared to the original form of the problem, the NNLS formulation does not explicitly promote sparsity by minimizing the 1-norm of the solution vector $\tilde{\bm{\rho}}^{v}$. Nevertheless, several numerical experiments—including the ones we report in this work—show that sparse solutions are obtained by the Lawson-Hanson algorithm [52, 62, 61].
@@
Before solving (A.1), we perform two conditioning operations on the constraints matrix $\bm{C}_{v}$, both to reduce the number of iterations required by the Lawson-Hanson algorithm to converge and to increase the derived solution’s accuracy [63]. First, we rescale each row of $\bm{C}_{v}$ so that its entries have absolute value less than or equal to one. Second, we use the reduced LQ decomposition of $\bm{C}_{v}=\bm{L}_{v}\bm{Q}_{v}$, where $\bm{L}_{v}\in\mathbb{R}^{N_{c}^{v}\times N_{c}^{v}}$ is lower triangular and $\bm{Q}_{v}\in\mathbb{R}^{N_{c}^{v}\times J_{v}}$ has orthonormal rows, to transform the NNLS problem (A.1) to
@@
where $\tilde{\bm{b}}=\bm{Q}_{v}\bm{\rho}^{v}\in\mathbb{R}^{N_{c}^{v}}$, under the assumption that $\bm{L}_{v}$ is nonsingular. In addition, we modify the error thresholds $\epsilon_{s}$ accordingly to ensure that the solution obtained from (A.3) will also satisfy (A.2) with the original error parameters. We define the new error thresholds $\tilde{\epsilon}_{s}$ by
@@
which is sufficient to ensure that the transformed problem’s solution satisfies (A.2) [63]. The transformed NNLS problem reads